# Entropic (Gromov) Wasserstein Flow Matching with GENOT

### v3

@@ 5.1 GENOT for Linear EOT problems
GENOT-L on a static SB benchmark. First, we show that GENOT accurately computes EOT couplings using a recent benchmark. For different $\varepsilon\in\{0.1,1,10\}$ and $d\in\{2,16,64,128\}$), Gushchin et al. (2023) consider pairs $\mu,\nu$ for which the $\pi^{\star}_{\varepsilon}$ solution of (LEOT) with $\ell_{2}^{2}$ cost $c(\mathbf{x},\mathbf{y})=\frac{1}{2}\|\mathbf{x}-\mathbf{y}\|_{2}^{2}$ is known. Tab. 1 shows that GENOT-L outperforms all baselines, choosing one configuration for all experiments as opposed to some baselines. Detailed results can be found in Tab. LABEL:table:bw_uvp and LABEL:table:cbw_uvp. Importantly, GENOT’s performance can be further improved when adapting the batch size, see Fig. LABEL:fig:eot_benchmark_all. These results are of particular significance since many baselines are tailored towards the specific task of solving (i) the linear EOT problem (LEOT), (ii) in the balanced setting (iii) with the $\ell_{2}^{2}$ cost. In contrast, the contributions of GENOT go beyond this restrictive setting. In the following, we demonstrate GENOT’s flexibility and highlight its relevance in various single-cell biology tasks.
U-GENOT-L on simulated data. To go beyond the balanced setting, Fig. 3 visualizes U-GENOT-L’s ability to learn unbalanced EOT plans and re-weighting functions. We further showcase the influence of $\tau_{1},\tau_{2}$ in Fig. LABEL:fig:1d_simulated_unbalanced_big.
@@
GENOT-L recovers single cell trajectories. We demonstrate GENOT-L’s capability to infer the development of cells on a dataset capturing gene expression of the developing mouse pancreas at embryonic days 14.5 (source) and 15.5 (target) (Bastidas-Ponce et al., 2019). As cells are known to evolve stochastically (Elowitz et al., 2002), we leverage GENOT’s capability to generate samples from the conditional distribution to model non-deterministic trajectories. Moreover, as Euclidean metrics often fail to represent meaningful distances in gene expression space (Moon et al., 2018), we leverage GENOT’s flexibility and use the geodesic distance on the data manifold $d_{\mathcal{M}}$ as cost, as discussed in § 3.1. Here, $\mathcal{M}=\mathcal{X}\cup\mathcal{Y}$ since the source and target domains have the same dimension. We also estimate the EOT coupling for the $\ell_{2}^{2}$ cost for comparison. Fig. 4 and Fig. LABEL:fig:pancreas_transitions highlight that the geodesic cost causes the conditional distribution to follow the data manifold. In particular, this helps to obtain biologically more plausible cell trajectories. We follow Gayoso et al. (2022) for assessing the uncertainty of cell trajectories by computing $\mathrm{cos}\text{-}\mathrm{var}(\pi_{\theta}(\cdot|\mathbf{x}))=\operatorname{Var}_{Y\sim\pi_{\theta}(\cdot|\mathbf{x})}[\mathrm{cos}\text{-}\mathrm{sim}(Y,\mathbb{E}_{Y\sim\pi_{\theta}(\cdot|\mathbf{x})}[Y])]$ See App. LABEL:app:metrics_general for more details on $\mathrm{cos}\text{-}\mathrm{var}$. We expect high uncertainty in cell types in early developmental stages and low variance in mature cell types. Indeed, Fig. 4 and Fig. LABEL:fig:pancreas_uncertainty_suppl show that the computed uncertainty is biologically meaningful: GENOT-L helps to uncover lineage branching events.
Figure 5: Accuracy of cellular response predictions of U-GENOT-L for different cancer drugs with varying unbalancedness parameter $\tau=\tau_{1}=\tau_{2}$. As $\tau$ decreases, we allow for more unbalancedness. For each $\tau$, we display results on 3 runs.
@@
U-GENOT-L predicts single-cell responses to perturbations. Neural OT has been successfully applied to model cellular responses to perturbations using deterministic maps (Bunne et al., 2021). GENOT has the comparative advantage to model conditional distributions, allowing for uncertainty quantification. We consider single cell RNA-seq data measuring the response of cells to 163 cancer drugs (Srivatsan et al., 2020). Each drug has been applied to a population of cells that can be partitioned into 3 different cell types. Source (resp. target) is cells before (resp. after) drug application. While there is no ground truth in the matching between unperturbed and perturbed cells due to the destructive nature of sequencing technologies, we know which unperturbed subset of cells is supposed to be mapped to which perturbed subset of cells. We use this to define an accuracy metric (see App. LABEL:app:metrics_cell_type_transition), while we assess the uncertainty of the prediction using $\mathrm{cos}\text{-}\mathrm{var}$. Fig. LABEL:fig:perturbation_calibration shows that for 117 out of 163 drugs, the model is perfectly calibrated (see App. LABEL:app:calibration_score), while it yields a negative correlation between error and uncertainty only for one drug. We can improve predictions by accounting for class imbalances between different cell types using U-GENOT. Indeed, Fig. 5 shows that allowing for mass variation increases accuracy for nine different cancer drugs, which we selected as they are known to have a strong effect. Fig. LABEL:fig:dacinostat and LABEL:fig:tanespimycin confirm the results visually.
@@
Figure 6: Mapping a Swiss roll in $\mathbb{R}^{3}$ ($\mu$, top left) to a spiral in $\mathbb{R}^{2}$ ($\nu$, bottom left) with GENOT-Q. Center: Color code tracks where samples from $\mu$ (top) are mapped to (bottom). Right column: samples from $\mu$ (top) and each of the corresponding conditionals, along with conditional density estimates. The learned QEOT coupling minimizes the cost distortion: Points close in support of $\mu$ are paired to points close in the support of $\nu$.
Figure 7: Top: UMAP of transported cells with GENOT-F (colored by cell type) and cells in the target distribution (gray). GENOT-F clusters similar cells together, i.e., according to their cell type. Bottom: UMAP of transported cells with a GENOT model trained on batch-wise independent couplings, thus not using OT. This remains to replace $\pi^{\star}_{\varepsilon}$ by $\mu\otimes\nu$ in Eq. (5). While the learned coupling then randomly mixes the cell types, it also degrades the global fitting of the target distribution of cells.
Figure 8: Benchmark of GENOT-Q models against discrete GW (GW-LR, App. LABEL:app:competing_methods) on translating cells between ATAC space of dimension $d_{1}$ and RNA space of dimension $d_{2}$. Performance is measured with the FOSCTTM score (App. LABEL:app:metrics_single_cell) and the Sinkhorn divergence between target and predicted target distribution. (left) intra-domain costs $c_{\mathcal{X}}=c_{\mathcal{Y}}=\ell_{2}^{2}$, (right) geodesic distances $c_{\mathcal{X}}=d_{\mathcal{X}}$ and $c_{\mathcal{Y}}=d_{\mathcal{Y}}$. Mean and 95% CI shown across 3 runs.
@@ 5.2 GENOT for Quadratic EOT Problems
GENOT-Q on simulated data. We transport a Swiss role in $\mathbb{R}^{3}$ to a spiral in $\mathbb{R}^{2}$. Fig. 6 shows that GENOT-Q successfully mimics an isometric alignment. Here, we set $\varepsilon=0.01$ and investigate its influence in more detail in Fig. LABEL:fig:gw_toy_data_eps.
@@
GENOT-Q translates modalities of single cells. The number of modalities of a cell that can be simultaneously measured is limited due to technical limitations. Yet, it is important to match measurements of different modalities to obtain a more holistic view of the profile of a cell. Demetci et al. (2022) address this task using discrete quadratic OT. However, due to the lack of neural quadratic EOT solvers, no OT-inspired method naturally allows for out-of-sample prediction. In the following, we use GENOT-Q to translate ATAC measurements (source) to gene expression space (target) on a bone marrow dataset (Luecken et al., 2021). As both modalities were measured in the same cell, the true match of each cell is known for this specific dataset. We compare GENOT-Q with discrete OT extended out-of-sample with linear regression (GW-LR, see LABEL:app:regression_discrete). We assess the performance using (i) the FOSCTTM (“Fractions of Samples Closer to the True Match”, see LABEL:app:foscttm) that measures the optimality of the learned coupling, and (ii) the Sinkhorn divergence (Feydy et al., 2019) between the predicted target and the target to assess distributional fitting. As in § 5.1, we leverage GENOT’s flexibility and use as intra-domains costs the geodesic distances on the source and target domain, namely $c_{\mathcal{X}}=d_{\mathcal{X}}$, $c_{\mathcal{Y}}=d_{\mathcal{Y}}$. We also estimate the EOT coupling for the $\ell_{2}^{2}$ cost for comparison. Results are shown in Fig. 8. Regarding the FOSCTTM score, we can see that (i) using geodesic costs is crucial in high dimensions and (ii) GW-LR is competitive in low dimensions but not for higher ones. Regarding distributional fitting, GENOT models outperform by orders of magnitude.
@@
GENOT-F improves modality translation. To enhance the performances attained by purely quadratic OT-based models, we introduce a novel method for translating between ATAC and RNA. We extend the idea proposed by Demetci et al. (2022) to the fused setting: We approximate RNA data from ATAC measurements using gene activity (Stuart et al., 2021), and we further process the data using a conditional VAE (Lopez et al., 2018a) to reduce batch effects. This way, we construct a joint space $\Omega$. Following the notations in § 3.3, RNA (source $\mathbf{x}\in\tilde{\mathcal{X}}$) and ATAC (target $\mathbf{y}\in\tilde{\mathcal{Y}}$) carry the structural information, while features $\mathbf{u}\in\Omega$ and $\mathbf{v}\in\Omega$ are obtained from the VAE embedding. Fig. LABEL:fig:fgw_modality_translation_diff_alphas shows that the additional linear penalty on the feature term helps to obtain a better alignment compared to GENOT-Q. Fig. 7 visualizes the learned fused coupling compared to GENOT trained with batch-wise independent couplings. When aligning multiple modalities of single cells, the proportions of cell types in the two modalities might differ. We simulate this setting by removing cells belonging to certain cell types Tab. LABEL:table:unbalancedness_ugenot_fgw shows that U-GENOT-F preserves high accuracy while learning meaningful rescaling functions.
Conclusion. We introduce GENOT, a versatile neural OT framework to learn cost-efficient stochastic maps, within the same space, or across incomparable spaces. We can loosen the mass conservation constraint if desired, while transport plans can be learned using any cost. We demonstrate that GENOT’s flexibility meets numerous requirements necessary for various prominent tasks in single-cell biology. GENOT models are implemented in ott-jax (Cuturi et al., 2022).
Impact Statement. This work introduces novel approaches in neural optimal transport and presents applications in single-cell biology. Leveraging single-cell genomics data holds promise for advancing personalized medicine but requires cautious handling due to its potential inclusion of sensitive information. Moreover, as neural optimal transport techniques find applicability across diverse domains, the societal ramifications of this research extend beyond single-cell genomics. With our publicly available implementation of GENOT in ott-jax (Cuturi et al., 2022), we enable its utilization by general machine learning practitioners. In a planned future release in moscot (Klein et al., 2023), we ensure GENOT’s accessibility to advance and accelerate research specifically tailored for single-cell biologists.
Acknowledgement. Co-funded by the European Union (ERC, DeepCell - 101054957). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. F.J.T. consults for Immunai Inc., Singularity Bio B.V., CytoReason Ltd, Cellarity, and has ownership interest in Dermagnostix GmbH and Cellarity.
@@
Appendix A Proofs
A.1 Proofs of § 1

### v4

@@ 5.1 GENOT-L for modeling cell trajectories
GENOT-L on simulated data We show that GENOT accurately computes EOT couplings using a recent benchmark [30]. Tab. LABEL:table:metrics_conditional_gaussian shows that GENOT-L outperforms all baselines considered. These results are of particular significance since most of the baselines are tailored towards the specific task of solving (i) the linear EOT problem (LEOT), (ii) in the balanced setting (iii) with the $\ell_{2}^{2}$ cost. In contrast, the contributions of GENOT go beyond this setting, see N2, N3, and N4. Going beyond the balanced problem (N3), we show that U-GENOT learns meaningful unbalanced EOT couplings between Gaussians [36] (LABEL:fig:unbalanced_eot), which we visualize in Fig. LABEL:fig:1d_simulated_unbalanced_big.
@@
GENOT-L learns the evolution of cells. Trajectory inference is a prominent task in single-cell biology. It is key to understand the development of cells from less mature to more mature cell states. In the following, we consider a dataset capturing gene expression of the developing mouse pancreas at embryonic days 14.5 (source) and 15.5 (target) [2]. Using the geodesic distance on the data manifold $d_{\mathcal{M}}$ as cost (§ 3.1) improves the learnt vector field compared to using the $\ell_{2}^{2}$ cost (TSI score in Fig.2, § LABEL:app:metrics_tsi). The CT error measures to what extent cells follow the manifold (§ LABEL:app:metrics_cell_type_transition, Fig. LABEL:fig:pancreas_transitions), which is visually confirmed by samples of the conditional distribution (Fig.2). These results support the need for a flexible choice of the cost function (N2). As cells are known to evolve stochastically, we leverage GENOT’s capability to generate samples from the conditional distribution to model non-deterministic trajectories (N1). We follow Gayoso et al. [25] for assessing the uncertainty of cell trajectories by computing $\mathrm{cos}\text{-}\mathrm{var}(\pi_{\theta}(\cdot|\mathbf{x}))=\operatorname{Var}_{Y\sim\pi_{\theta}(\cdot|\mathbf{x})}[\mathrm{cos}\text{-}\mathrm{sim}(Y,\mathbb{E}_{Y\sim\pi_{\theta}(\cdot|\mathbf{x})}[Y])]$ (§ LABEL:app:metrics_general). We expect high uncertainty in cell types in early developmental stages and low variance in mature cell types. Indeed, Fig. 2 and Fig. LABEL:fig:pancreas_uncertainty_suppl show that the computed uncertainty is biologically meaningful.
@@
U-GENOT-L predicts single-cell responses to perturbations. Neural OT has been successfully applied to model cellular responses to perturbations using deterministic maps [7]. GENOT has the comparative advantage to model conditional distributions, allowing for uncertainty quantification. We consider single cell RNA-seq data measuring the response of cells to 163 cancer drugs [78]. Each drug has been applied to a population of cells that can be partitioned into 3 different cell types. The source (resp. target) distribution consists of cells before (resp. after) drug application. While there is no ground truth in the matching between unperturbed and perturbed cells due to the destructive nature of sequencing technologies (hence the need for realignment), we know which unperturbed subset of cells is supposed to be mapped to which perturbed subset of cells. We use this to define an accuracy metric (see App. LABEL:app:metrics_cell_type_transition), while we assess the uncertainty (N1) of the prediction using $\mathrm{cos}\text{-}\mathrm{var}$. Fig. LABEL:fig:perturbation_calibration shows that for 117 out of 163 drugs, the model is perfectly calibrated (see App. LABEL:app:calibration_score), while it yields a negative correlation between error and uncertainty only for one drug. We can improve the generation of cells by accounting for class imbalances between different cell types using U-GENOT. These imbalances might occur due to biases in the experimental setup or due to cell death [78]. Indeed, Fig. 3 shows that allowing for mass variation increases accuracy for nine different cancer drugs, selected as they are known to have a strong effect. Fig. LABEL:fig:dacinostat and LABEL:fig:tanespimycin confirm the results visually (N3).
@@
Figure 3: Left: Accuracy of cellular response predictions of U-GENOT-L for cancer drugs with varying unbalancedness parameter $\tau=\tau_{1}=\tau_{2}$. Smaller $\tau$ implies more unbalancedness (3 runs per $\tau$). Right: Mapping a Swiss roll in $\mathbb{R}^{3}$ ($\mu$) to a spiral in $\mathbb{R}^{2}$ ($\nu$) with GENOT-Q. Center: Color code tracks where samples from $\mu$ (top) are mapped to (bottom). Right column: samples from $\mu$ (top) and the corresponding conditionals, along with conditional density estimates. The learned QEOT coupling minimizes the distortion: Points close in support of $\mu$ generate points close in support of $\nu$.
@@ 5.2 GENOT for Quadratic EOT Problems
As outlined in § 1, quadratic OT problems are of particular interest in single-cell genomics (N4). Here, we focus on the task of aligning different cellular modalities, which is necessary as most sequencing technologies are limited to measuring one modality [34]. Yet, there exist few protocols measuring two modalities simultaneously, thus allowing us to assess the alignment as the ground truth matching is known. We leverage this to assess GENOT-Q’s performance on the task of translating cells from ATAC-seq to RNA-seq, and also propose a novel way to improve the alignment using GENOT-F. Before, we demonstrate the task of mapping across incomparable spaces on simulated data.
@@
GENOT-Q and U-GENOT-Q on simulated data. To visualize the problem of quadratic OT, we transport a Swiss role in $\mathbb{R}^{3}$ to a spiral in $\mathbb{R}^{2}$. Fig. 3 shows that GENOT-Q successfully mimics an isometric alignment. Here, we set $\varepsilon=0.01$ and investigate its influence in more detail in Fig. LABEL:fig:gw_toy_data_eps. To quantitatively assess the performance of GENOT-Q and U-GENOT-Q, we compute (unbalanced) entropic OT plans between Gaussian distributions, which the closed form EOT plan is known for [43]. Fig.LABEL:fig:gromov_eot_gaussian and Fig.LABEL:fig:unbalanced_gw_eot_gaussians show that GENOT-Q is able to meaningfully learn (un-)balanced entropic EOT plans for lower dimensions, but its performance decreases with increasing dimensionality.
@@
GENOT-Q translates modalities of single cells. We use GENOT-Q to translate ATAC measurements (source) to gene expression space (target) on a bone marrow dataset [51]. As both modalities were measured in the same cell, the true match of each cell is known for this specific dataset. We compare GENOT-Q with discrete OT extended out-of-sample with linear regression (GW-LR, see LABEL:app:regression_discrete). We assess the performance using (i) the FOSCTTM (“Fractions of Samples Closer to the True Match”, see LABEL:app:foscttm) that measures the optimality of the learned coupling, and (ii) the Sinkhorn divergence [23] between the predicted target and the target to assess distributional fitting. As in § 5.1, we leverage GENOT’s flexibility and use as intra-domains costs the geodesic distances on the source and target domain, namely $c_{\mathcal{X}}=d_{\mathcal{X}}$, $c_{\mathcal{Y}}=d_{\mathcal{Y}}$ (N2). We also estimate the EOT coupling for the $\ell_{2}^{2}$ cost for comparison. Results are shown in Fig. 3. Regarding the FOSCTTM score, we see that (i) using geodesic costs is crucial in high dimensions and (ii) GW-LR is competitive in low dimensions but not for higher ones. Regarding distributional fitting, GENOT models outperform by orders of magnitude.
Figure 4: Left: Benchmark of GENOT-Q models against discrete GW (GW-LR, App. LABEL:app:competing_methods) on translating cells between ATAC space of dim. $d_{1}$ and RNA space of dim. $d_{2}$, with performance measured with FOSCTTM score (App. LABEL:app:metrics_single_cell) and Sinkhorn divergence between target and predicted target distribution. (left) intra-domain costs $c_{\mathcal{X}}=c_{\mathcal{Y}}=\ell_{2}^{2}$, (right) geodesic distances $c_{\mathcal{X}}=d_{\mathcal{X}}$ and $c_{\mathcal{Y}}=d_{\mathcal{Y}}$. We show mean and std across 3 runs. Right: Top: UMAP of transported cells with GENOT-F (colored by cell type) and cells in the target distribution (gray). Cells of the same cell type generate cells which cluster together in RNA space. Bottom: UMAP of transported cells with a GENOT model trained on batch-wise independent couplings, thus not using OT, generating cells which are randomly mixed.
@@
GENOT-F improves modality translation. To enhance the performances attained by purely quadratic OT-based models, we introduce a novel method for translating between ATAC and RNA. We extend the idea of translating between cell modalitiies proposed by Demetci et al. [17] to the fused setting: We approximate RNA data from ATAC measurements using gene activity [79], and we further process the data using a conditional VAE [48] to reduce batch effects. This way, we construct a joint space $\Omega$. Following the notations in § 3.3, RNA (source $\mathbf{x}\in\tilde{\mathcal{X}}$) and ATAC (target $\mathbf{y}\in\tilde{\mathcal{Y}}$) carry the structural information, while features $\mathbf{u}\in\Omega$ and $\mathbf{v}\in\Omega$ are obtained from the VAE embedding. Fig. LABEL:fig:fgw_modality_translation_diff_alphas shows that the additional linear penalty on the feature term helps to obtain a better alignment compared to GENOT-Q. Fig. 4 visualizes the learned fused coupling compared to GENOT trained with batch-wise independent couplings. When aligning multiple cell modalities, the proportions of cell types in the two modalities often differ (N3). We simulate this setting by removing cells belonging to certain cell types. Tab. LABEL:table:unbalancedness_ugenot_fgw shows that, in this unbalanced setting, U-GENOT-F preserves high accuracy while learning meaningful rescaling functions.
Conclusion. Motivated by emerging applications in single-cell genomics, we introduce GENOT, a versatile neural OT framework to learn cost-efficient stochastic mappings, both within the same space or across incomparable spaces. GENOT can be effectively used for learning EOT couplings with any cost function, and flexibly allows for loosening mass constraints if desired.
@@
Appendix A Limitations
While our study opens up the use of neural optimal transport to a wide range of tasks in single-cell genomics, there are limitations of our work:
  • The solution to the entropic (Fused) Gromov-Wasserstein problem is not necessarily unique. As GENOT-Q and GENOT-F leverage discrete solvers in each iteration, the orientation of the solution across iterations is not maintained. To prevent learning a mixture of solutions, we propose to use a fused term whenever possible, make use of different cost functions if helpful, or using GENOT with our proposed initialisation scheme (§ LABEL:app:genot_quadratic). While the latter can always be applied, its major drawback is the costly training as a neural ODE has to be solved at every iteration.
  • When learning unbalanced EOT plans, choosing a value for the hyperparameters $(\tau_{a},\tau_{b})$ is not evident. As discussed in Séjourné et al. [75], the choice of these parameters is mostly empirical.
  • While GENOT is motivated by the extensive use of optimal transport in single-cell genomics, there is no evidence that cells evolve exactly according to an entropic OT plan.
  • Evaluations of learnt transport plans for tasks in single-cell genomics experiments are task-specific, and do not necessarily reflect to what extent the learnt coupling is an entropic OT coupling. In effect, we rely on prior biological knowledge (CT error, Fig. 2) or on common downstream trajectory inference methods in single-cell genomics to evaluate the learnt velocity field (TSI score, Fig. 2) when computing EOT plans for trajectory inference. Similarly, we can only assess learnt EOT plan on a cluster level, and not a single data point level when studying cellular perturbations (3). For the task of modality translation (Fig. 4), we know the ground-truth coupling. Yet, the ground-truth coupling is deterministic and one-to-one, and hence is not a valid solution of an entropy-regularized Gromov-Wasserstein OT plan with $\varepsilon>0$.
Appendix B Propositions and Proofs
B.1 Existence of EOT couplings
In this section, we point out that a solution to the entropic fused GW problem  (EFGW), and its unbalanced counterpart, always exists. Note that for $\alpha=0$, we recover the existence for the pure entropic GW (EFGW) and its unbalanced counterpart (UQEOT) problems. The arguments are adapted from [88, Prop. 2]
Balanced setting: We first recall the fused Gromov-Wasserstein problem below.
$\min_{\pi\in\Pi(\mu,\nu)}\int_{((\Omega\times\tilde{\mathcal{X}})\times(\Omega\times\tilde{\mathcal{Y}}))^{2}}D_{c_{\tilde{\mathcal{X}}},c_{\tilde{\mathcal{Y}}}}^{c}\mathrm{d}(\pi\otimes\pi)+\varepsilon\mathrm{KL}(\pi|\mu\otimes\nu)$ (6)
where $D_{c_{\tilde{\mathcal{X}}},c_{\tilde{\mathcal{Y}}}}^{c}\left((\mathbf{u},\mathbf{x}),(\mathbf{v},\mathbf{y}),\mathbf{x}^{\prime},\mathbf{y}^{\prime}\right):=(1-\alpha)c(\mathbf{u},\mathbf{v})+\alpha|c_{\tilde{\mathcal{X}}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{\tilde{Y}}}(\mathbf{y},\mathbf{y}^{\prime})|^{2}$. We first need to show that a solution $\pi^{\star}_{\varepsilon}$ to this problem always exists. We assume that the intra-domain $c_{\mathcal{X}}$, $c_{\mathcal{Y}}$ costs, and the inter-domain cost $c$ are l.s.c, an assumption that matches all our experimental settings. From [88, Lemma 3], this implies that the functional:
$\pi\in\Pi(\mu,\nu)\mapsto\int_{((\Omega\times\tilde{\mathcal{X}})\times(\Omega\times\tilde{\mathcal{Y}}))^{2}}D_{c_{\tilde{\mathcal{X}}},c_{\tilde{\mathcal{Y}}}}^{c}\mathrm{d}(\pi\otimes\pi)$ (7)
is weakly l.s.c, i.e. l.s.c for the convergence in law. Moreover, from [23, Prop. 8], $\pi\mapsto\mathrm{KL}(\pi|\mu\otimes\nu)$ is also weakly l.s.c. Therefore, the objective function of the fused Gromov-Wasserstein problem:
$\mathcal{F}:\pi\in\Pi(\mu,\nu)\mapsto\int_{((\Omega\times\tilde{\mathcal{X}})\times(\Omega\times\tilde{\mathcal{Y}}))^{2}}D_{c_{\tilde{\mathcal{X}}},c_{\tilde{\mathcal{Y}}}}^{c}\mathrm{d}(\pi\otimes\pi)+\varepsilon\mathrm{KL}(\pi|\mu\otimes\nu)$ (8)
is weakly l.s.c, as a sum of weakly l.s.c functionals. Then, since we assume that $\mathcal{X}$ and $\mathcal{Y}$ are compact, they are polish spaces, so from [68, Theorem 1.7], $\Pi(\mu,\nu)$ is weakly compact. Therefore, from Weierstrass Theorem, $\mathcal{F}$ attains its minimum on $\Pi(\mu,\nu)$, i.e. it exists $\pi^{\star}_{\varepsilon}$ solution to this problem.
Unbalanced setting: The proof is very similar. We first recall the unbalanced fused Gromov-Wasserstein problem.
$\min_{\pi\in\mathcal{M}^{+}(\mathcal{X}\times\mathcal{Y})}\int_{((\Omega\times\tilde{\mathcal{X}})\times(\Omega\times\tilde{\mathcal{Y}}))^{2}}D_{c_{\tilde{\mathcal{X}}},c_{\tilde{\mathcal{Y}}}}^{c}\mathrm{d}(\pi\otimes\pi)+\varepsilon\mathrm{KL}(\pi|\mu\otimes\nu)+\lambda_{1}\mathrm{KL}^{\otimes}(\pi_{1}|\mu)+\lambda_{2}\mathrm{KL}^{\otimes}(\pi_{2}|\mu)$ (9)
where $\mathrm{KL}^{\otimes}(\alpha|\beta)=\mathrm{KL}(\alpha\otimes\alpha|\beta\otimes\beta)$. From [81, Prop. 8], $\mathrm{KL}^{\otimes}$ is weakly l.s.c, so is $\pi\mapsto\mathrm{KL}^{\otimes}(\pi_{1}|\mu)$ and $\pi\mapsto\mathrm{KL}^{\otimes}(\pi_{2}|\nu)$. Therefore, the objective function of the unbalanced fused Gromov-Wasserstein problem is also weakly l.s.c. as a sum of weakly l.s.c. functions. Moreover, since $\mathcal{X}$ and $\mathcal{Y}$ are assumed to be compact, $\mathcal{X}\times\mathcal{Y}$ is also compact and the Banach-Alaoglu provides the compactness of $\mathcal{M}^{+}(\mathcal{X}\times\mathcal{Y})$. Then, from Weierstrass Theorem, similarly to the balanced setting, we get the existence of a solution $\pi^{\star}_{\varepsilon,\tau}$ to this problem.
B.2 Proofs of § 3.2